10 ml of boric acid solution was placed inside 250ml conical flask And after adding distilled water and mannitol to the solution it required 15ml of .10 M naoh solution to reach phenolphthalein endpoint the molarity of the boric acid solution is


1
Expert's answer
2021-06-11T22:48:52-0400

Molarity "= moles\/ litre"

Moles=15/61.8

=0.2427moles

Molarity = 0.2427/1000

= 2.42× 10^-4M